# SeamPose: Repurposing Seams as Capacitive Sensors in a Shirt for Upper-Body Pose Tracking

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; San Francisco, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 4.3. SeamPose Signals
Changes in the wearer’s upper body pose cause changes in the seam electrode’s sensed capacitances, which are input into our deep learning pipeline. When connected to the sensing board, the conductive seams become self-capacitance sensor electrodes. In a loading-mode capacitive sensing system, the setup is simple: each sensor electrode acts as both a transmitter and a receiver (Smith, 1996). In any loading-mode system, electrical current displacement occurs as the electrode’s capacitive coupling with the surrounding changes. In SeamPose, the seam electrodes are coupled with the wearer’s body, so when the body pose changes, leading to changes in the coupling between the body and seam electrodes, hence the current displacement changes. In addition, the seam electrode itself deforms, distorts, and displaces during body movements, altering the electrode’s capacitance which also affects the current displacement. In a nutshell, we conjecture the capacitive changes are attributed to (1) coupling changes between the seam electrodes and the wearer’s body, and (2) seam electrode’s self-capacitance changes.
@@
As shown in Fig. 5, the seam signals correlate to the wearer’s movements. The sleeve electrodes, colored purple, have the largest changes in magnitude as they are the longest electrodes with the largest range of motion. The shoulder front (colored beige) and back (colored green) electrodes have similar magnitudes as they are symmetrically placed, but the relationship between these two signals depends on the motion. For example, for the both-arm side raises, the two signals both increase as the arms raise and decrease as the arms return, but for alternate punching movements, one signal increases as the other decreases. The shoulder top (colored blue) has the smallest magnitude for its shortest length and relatively stationary nature, compared with other electrodes that touch moving joints. Though the magnitude is small, it is proven to still be information-rich in our later evaluation. Although our prototype does not cover or directly instrument the head, it can sense the head movement (see head roll signals differing from standing still signals in Fig. 5) based on overall body-capacitance changes and subtle shoulder movements. To extract and interpret the pose information from these complex capacitance changes, we developed a data-driven approach.
@@ 4.4. Deep Learning Pipeline
Figure 6. Customized Deep Learning Pipeline.
With eight channels of capacitive signals from the seams on the shirt, we estimate the wearer’s upper body pose in 3D (x,y,z coordinates of 8 joints), relative to the pelvis. In the previous subsection, we show the correlation between body movements and seam signals, but the reconstruction task is challenging because the tracked body parts have a total of 19 degrees of freedom (Cazzola et al., 2017), and we only have sparsed 8 channels of 1-dimensional temporal signals, constrained by our design choice that minimally alters everyday clothing. Unlike tight-fitting suits, the sensors are not firmly fixed to a mapped body part, complicating the problem (Zuo et al., 2024a). The task is further complicated by the soft cotton fabric having natural draping variations. For example, when the wearers raise their arms over the head, the sleeves naturally slide down the arms and attribute signal variations. Embracing these technical hurdles in pursuit of comfortable and seamless pose-tracking integration into everyday clothing, we customized a deep-learning pipeline.
@@ 4.4.1. Ground Truth Acquisition
Recent computer vision advancements in pose reconstruction with a single RGB image have proven more accurate for users in loose-fitting cloth (Ray et al., 2023). Similar to prior works on non-vision-based pose tracking (Chen et al., 2024; Zhou et al., 2023), we selected state-of-the-art computer vision models for ground truth acquisition. Given an RGB image frame, Detectron2777https://github.com/facebookresearch/detectron2 first detects the area containing the person, then HMR2.0 (Goel et al., 2023), a vision transformer-based model, estimates the SMPL (Loper et al., 2015) representation of the detected person. SMPL is a mathematical model that describes a human body mesh template of 6890 vertices with shape components $\beta$ (dim($\beta$)=10) and pose components $\theta$ (dim($\theta$)= 216 = 24 joints x (3 x 3) rotation matrix representing the rotation from its kinematic prior). Among the 24 joint components, the pelvis provides global orientation (i.e. the direction the body is facing) and the other 23 joint components represent rotations from their kinematic prior. SMPL also acts as a forward-kinematic model and calculates joint positions based on joint rotations. Note, SeamPose does not predict body shape or global orientation. The ground truth labels have 2 folds:
@@ 4.4.3. Model
The model architecture is detailed in Fig. 6. The input signals, of dimension [96x8], are first transformed into an embedding of dimension 96 with a linear layer. Then, the embeddings go through a 2-layer bidirectional LSTM (inspired by (Mollyn et al., 2023; Chen et al., 2023; Huang et al., 2018; Zuo et al., 2024b)) of hidden dimension 256. We take the last 3 prediction frames (0.09s) and decode the embeddings with 2 linear layers with out-feature sizes of 256 and 78 (for the 13 pose components). We then MedianPool the last 3 prediction frames to mitigate jitter (Chen et al., 2023). Finally, we use SMPL to calculate joint positions. We further smooth the predictions with a running median filter of window size 5 (0.17s).
@@ 5. User Study Evaluation
For each participant, we collect 8 sessions of data. Before each session, the participant was instructed to take off and put back on the shirt themself in order to evaluate our system across different wearing sessions. Participants wore their own clothes of various types and sizes that fit underneath our prototype shirt. Each session lasted about 227 seconds. Each session contains 3 sections, including three types of upper-body movements informed by prior work:
@@ 6.2. Unseen User: User-Independent Model Results
For an even better user experience, the shirt should be able to track body poses "out-of-box" without any calibration data from the new user, i.e. user-specific training data. We adopt a leave-one-participant (LOPO) cross-validation to evaluate our performance in this scenario. User-independent (UI) models were trained 7.75 hours of data: 30.2 minutes of data from each of the other 11 participants and 5 hours of data from 7 researchers. As shown in Fig. 9, the MPJPE increases to 8.6cm (std=0.93cm), 1.6cm worse than 3.8min of training data, and 2.6cm worse than 22.7min of training data. The performance degradation is expected because individuals’ body sizes and capacitances vary, but SeamPose still yields low-fidelity tracking, sufficient for certain applications like activity recognition and lifelogging.
@@ 6.4. Results Analysis based on Motion
We further break down the results into motion types. Like many other data-driven wearable MoCap systems, the performance decreased for estimating unseen gestures/movements. As we discussed in Sec. 5, we cannot exhaust all possible upper-body movements&poses for evaluation. However, when we designed the study, we explicitly included dance and freestyle movements to better understand our system’s limitations on extreme and unseen poses and movement patterns. As shown in Fig. 10, across all three models, defined movements (5.9cm, 6.8cm, and 8.3cm with 22.7min, 3.8min, and 0min of user data) have the smallest errors, followed by dance movements (8.2cm, 9.5cm, and 10.8cm with 22.7min, 3.8min, and 0min of user data) and freestyle movements (10.8cm, 12.1cm, and 13.5cm with 22.7min, 3.8min, and 0min of user data). Adding user-specific training data also consistently improves the tracking accuracy for each movement type. This result suggested that the system similar to many data-driven systems, can further benefit from being trained with a much larger dataset containing a large variety of body movements. However, as a proof of concept, it still showed reliable tracking performance on unseen movements and subjects.
@@ 4 IMUs
As we discussed in Sec. 4.3 the shoulder top seams have the smallest changes in magnitude, removing them does not harm the performance greatly: 0.2cm, 0.2cm, and 0.5cm worse with 22.7min, 3.8min, and 0min of user data. The shoulder front/back seams have similar significance across all models, and removing them increases the errors by 1cm/0.9cm, 0.8cm/1cm, and 1cm/1.1cm with 22.7min, 3.8min, and 0min of user data. On the other hand, the sleeve electrodes’ importance varies across models. For both adaptive models, the sleeve electrodes have the most important placements (i.e. their removal causes the largest errors compared with other placements) which lead to about 1cm additional error when removed. For the user-independent model, the sleeve electrodes become the least important placement and only lead to an additional error of 0.3cm. The observation implies that signals from some placements (sleeve) generalize across users worse than others do. Furthermore, given the reasonable performance across all three models with the sleeve electrodes removed, SeamPose has the potential to generalize on short-sleeve shirts and even sleeveless ones.
6.6. Performance Comparison with Prior Works

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676341isbn: 979-8-4007-0628-8/24/10
@@
[…]
[]This figure shows the overall system overview of SeamPose. On the left, it shows colorful insulated conductive thread sewn over conventional seams on a black long-sleeve t-shirt. There are 4 different colors used for the thread: purple, blue, green, beige, with each representing a different channel and symmetric sensing channels being the same color. All the seams meet at a device on the upper back portion of the shirt which sends signals wirelessly to a nearby computer for processing. On the right side of the figure, it shows a SeamPose user in a pose along with a glimpse of the signals sent by SeamPose fed into a DL model that creates a 3D upper-body pose estimation of the user.
@@ 3. Seams & Patterns
[…]
[]This figure is a simplified view of the four black fabric pieces: left sleeve, front panel, back panel, and right sleeve. These pieces compose the pattern of a long-sleeve black t-shirt used and the figure indicates with colored dotted lines, which match the sensing channel colors, where the conductive thread is sewn on the fabric over the existing seams.
@@ 4.1. Conductive Seams Fabrication
[…]
[]This figure is composed of 4 subfigures: A, B, C, and D. A shows the home sewing machine setup. B is an image of the off-the-shelf insulated conductive thread used by SeamPose. C shows an illustration of how the machine-sewn conductive thread is traced over existing seams on the t-shirt. D shows a side-view illustration of how the TPU-insulated conductive thread with a silver nylon core is stitched onto the fabric. In both C and D the insulating TPU is highlighted in blue, the polyester sewing thread is purple, and the conductive silver nylon is dark blue.
@@
[…]
[]This figure has two parts, on the left side it shows a picture of the back of a user wearing SeamPose. From the outside, SeamPose looks like a normal black long-sleeve t-shirt, except for the 3D-printed customized sensing board housing hot-glued below the neck. On the right side of the figure, there are two images of the customized sensing board. The top image shows the front of the sensing board and a length measurement of 36mm. The bottom image shows the back of the sensing board and a width measurement of 31mm.
@@
[…]
[]On the left side of this figure, it shows a drawing of SeamPose with the sensing electrodes highlighted and color-coded: Sleeve is purple, Shoulder Front is beige, Shoulder Back is green, Shoulder Top is blue. On the top right side, there are images of a user on top performing various movements with graphs of the Left and Right seams at the bottom displaying the corresponding SeamPose signals. The colors of the signals correlate with the sensing channels.
@@ 4.3. SeamPose Signals
Changes in the wearer’s upper body pose cause changes in the seam electrode’s sensed capacitances, which are input into our deep learning pipeline that infers the relationship between the body pose and measurement capacitances. When connected to the sensing board, the conductive seams become self-capacitance sensor electrodes. In a loading-mode capacitive sensing system, the setup is simple: each sensor electrode acts as both a transmitter and a receiver (Smith, 1996). In any loading-mode system, electrical current displacement occurs as the electrode’s capacitive coupling with the surrounding changes. In SeamPose, the seam electrodes are coupled with the wearer’s body, so when the body pose changes (i.e., limb movements), the seams move along with the limbs, leading to changes in the coupling between the body and seam electrodes, hence the current displacement changes. In addition, the seam electrode itself deforms, distorts, and displaces during body movements, altering the electrode’s capacitance which also affects the current displacement. In a nutshell, we conjecture the capacitive changes are attributed to (1) coupling changes between the seam electrodes and the wearer’s body, and (2) seam electrode’s self-capacitance changes.
@@
As shown in Fig. 5, the seam signals correlate to the wearer’s movements. The sleeve electrodes, colored purple, have the largest changes in magnitude as they are the longest electrodes with the largest range of motion. The shoulder front (colored beige) and back (colored green) electrodes have similar magnitudes as they are symmetrically placed, but the relationship between these two signals depends on the motion. For example, for the both-arm side raises, the two signals both increase as the arms raise and decrease as the arms return, but for alternate punching movements, one signal increases as the other decreases. The shoulder top (colored blue) has the smallest magnitude for its shortest length and relatively stationary nature, compared with other electrodes that touch moving joints. Though the magnitude is small, it is proven to still be information-rich in our later evaluation.
Figure 6. Customized Deep Learning Pipeline.
[…]
[]The figure shows the Customized Deep Learning Pipeline used by SeamPose. The raw 8-channel signals are put into a normalized input window with 96 frames, equivalent to 3 seconds. This is then fed into an FC layer with a dropout of 0.2, activated by LeakyReLU. Then the embeddings go through a 2-layer Bi-LSTM with a dropout of 0.3 and hidden size of 256. The last 3 frames, out of 96 frames, of the output 512 embeddings are then fed into a fully connected later layer with a dropout of 0.2, LeakyReLU activated. Then the embeddings go through a final fully-connected layer, activated by Tahn. The output of dimension 3x78 goes through medianPool into dimension 78 and corresponds to 13x6 pose components. Finally based on the 13x6 pose components, SMPL calculates 8x3 joint positions.
For asymmetric movements (e.g., right arms curls and left front raises), the static side still exhibits signal changes with small magnitudes. Moving one side of the shirt often stretches the other side because different parts of the shirt are interconnected and body movements involve complex muscle engagements. For example, a one-arm raise could (a) stretch the back of the shirt which changes the signals as the sensing board is in the back, and (b) alter the whole body’s coupling with the shirt. Further, at the thread-to-sensing-board connection, some insulated threads are close to and even overlap with each other and thus attributed to crosstalks, which is a limitation to be addressed in the future, but in our prototype, we input the 8 seam signals together, instead of side-by-side, to avoid possible confusion to the DL model.
Although our prototype does not cover or directly instrument the head, it can sense the head movement (see head roll signals differing from standing still signals in Fig. 5) based on overall body-capacitance changes and subtle shoulder movements. To extract and interpret the pose information from these complex capacitance changes, we developed a data-driven approach.
@@ 4.4. Deep Learning Pipeline
With eight channels of capacitive signals from the seams on the shirt, we estimate the wearer’s upper body pose in 3D, relative to the pelvis. In the previous subsection, we show the correlation between body movements and seam signals, but the complex mapping between the 8 channels and the body pose is not immediately clear. The reconstruction task is challenging because the tracked body parts have a total of 19 degrees of freedom (Cazzola et al., 2017), and we only have sparsed 8 channels of 1-dimensional temporal signals, constrained by our design choice that minimally alters everyday clothing. Unlike tight-fitting suits, the sensors are not firmly fixed to a mapped body part, complicating the problem (Zuo et al., 2024a). The task is further complicated by the soft cotton fabric having natural draping variations. For example, when the wearers raise their arms over the head, the sleeves naturally slide down the arms and attribute signal variations. Embracing these technical hurdles in pursuit of comfortable and seamless pose-tracking integration into everyday clothing, we customized a deep-learning pipeline to learn the complex mapping between the 8 input channels.
@@ 4.4.1. Ground Truth Acquisition
Recent computer vision advancements in pose reconstruction with a single RGB image have proven more accurate for users in loose-fitting cloth (Ray et al., 2023). Similar to prior works on non-vision-based pose tracking (Chen et al., 2024; Zhou et al., 2023; Li et al., 2024, 2022; Lee et al., 2024; Yu et al., 2024), we selected state-of-the-art computer vision models for ground truth acquisition. Given an RGB image frame, Detectron2777https://github.com/facebookresearch/detectron2 first detects the area containing the person, then HMR2.0 (Goel et al., 2023), a vision transformer-based model, estimates the SMPL (Loper et al., 2015) representation of the detected person. SMPL is a mathematical model that describes a human body mesh template of 6890 vertices with shape components $\beta$ (dim($\beta$)=10) and pose components $\theta$ (dim($\theta$)= 216 = 24 joints x (3 x 3) rotation matrix representing the rotation from its kinematic prior). Among the 24 joint components, the pelvis provides global orientation (i.e., the direction the body is facing) and the other 23 joint components represent rotations from their kinematic prior. SMPL also acts as a forward-kinematic model and calculates joint positions, $J$, based on joint rotations, $\theta$. Note, SeamPose does not predict body shape or global orientation. The ground truth labels have 2 folds:
@@ 4.4.3. Model
The model architecture is detailed in Fig. 6. The input signals, of dimension [96x8], are first transformed into an embedding of dimension 96 with a linear layer. Then, the embeddings go through a 2-layer bidirectional LSTM (inspired by (Mollyn et al., 2023; Chen et al., 2023; Huang et al., 2018; Zuo et al., 2024b)) of hidden dimension 256. LSTM is chosen to better model the temporal patterns, but there are potential optimizations on model architectures (e.g., transformers) for future works. We take the last 3 prediction frames (0.09s) and decode the embeddings with 2 linear layers with out-feature sizes of 256 and 78 (for the 13 pose components). We then MedianPool the last 3 prediction frames to mitigate jitter (Chen et al., 2023). Finally, we use SMPL to calculate joint positions based on the pose components. Our model directly regresses the pose components representing joint angles that are more transferrable across body sizes. The conversion to joint positions allows easy comparisons with other systems as the position error is the common metric for wearable pose tracking. We further smooth the joint position predictions with a running median filter of window size 5 (0.17s).
@@ 4.5. Real-Time Inference Pipeline
[…]
[]This figure shows four example screenshots of real-Time inference with SeamPose: with the wearer wearing SeamPose on the left and predictions on a laptop screen on the right. In the first example, the user is learning to their left. In the second example, the user is T-posing. In the third example, the user is doing a bicep curl with their left arm. In the fourth example, the user is in a boxing stance.
@@ 5. User Study Evaluation
For each participant, we collected 8 sessions of data. Before each session, the participant was instructed to take off and put back on the shirt themself in order to evaluate our system across different wearing sessions. Participants wore their own clothes of various types and sizes that fit underneath our prototype shirt. Each session lasted about 227 seconds. Each session contains 3 sections, including three types of upper-body movements informed by prior work:
[…]
[]This is composed of 2 subfigures: A and B. A shows 3 color-coded bar graphs with the distance error distributions of: a User-Adaptive Model with 22.68 minutes of User Data, a User-Adaptive Model with 3.78 minutes of User Data, and a User-Independent Model with 0 minutes of User Data. The Y-axis of each graph is the Euclidean Distance Error in centimeters. The X-axis of each graph has the data points for: Head, Neck, Left Shoulder, Left Elbow, Left Wrist, Right Shoulder, Right Elbow, Right Wrist. For the user adaptive model with 22.7 min of training data, the user adaptive model with 3.78 min of training data, and the user-independent model with 0 min of user data, respectively, the head errors are 3.3, 3.8, and 6.4 cm; the neck errors are 2.0, 2.3, and 4.0 cm; the left shoulder errors are 2.3, 2.6, and 3.7cm; the left elbow errors are 6.6, 7.6, and 8.9cm; the left wrist errors are 12.6, 14.3, and 16.5cm; the right shoulder errors are 2.3, 2.6, and 3.7cm, the right elbow errors are 6.9, 8.0, and 9.3cm; the right wrist errors are 12.5, 14.6, and 16.4cm. B shows a stick-figure representation of a user’s upper body with each of the joints labeled and color-coded, corresponding to the labels in (A).
@@ 6. Results
[…]
[]This figure shows a line graph that shows how the model’s tracking performance increases when more training data. The y-axis is the Mean Per Joint Position Error in centimeters and the x-axis is the amount of training data in minutes at a increment of 3.7min: the error is 8.6cm 0 minutes of training data; the error is 7.0 cm with 3.7min of training data; the error is 6.5cm with 7.4 min of training; the error is 6.4cm with 11.1min of training data; the error is 6.2 with 14.8min of training data; the error is 6.1cm with 18.5min of training data; the error is 6.0 with 22.7min of training data.
@@ 6.2. Unseen User: User-Independent Model Results
For an even better user experience, the shirt should be able to track body poses "out-of-box" without any calibration data from the new user, i.e., user-specific training data. We adopt a leave-one-participant (LOPO) cross-validation to evaluate our performance in this scenario. User-independent (UI) models were trained 7.75 hours of data: 30.2 minutes of data from each of the other 11 participants and 5 hours of data from 7 researchers. As shown in Fig. 9, the MPJPE increases to 8.6cm (std=0.93cm), 1.6cm worse than 3.8min of training data, and 2.6cm worse than 22.7min of training data. The performance degradation is expected because individuals’ body sizes and capacitances vary, but SeamPose still yields low-fidelity tracking, sufficient for certain applications like activity recognition and lifelogging.
@@ 6.3. Results Analysis based on Joints
[…]
[]This figure shows 3 bar graphs with movement error break of: a User-Adaptive Model with 22.68 minutes of User Data, a User-Adaptive Model with 3.78 minutes of User Data, and a User-Independent Model with 0 minutes of User Data. The y-axis is the Mean Per Joint Position Error (MPJPE). The bar graph bars are color-coded with purple being the 54 defined movements, green being the dance movements, and blue being the freestyle movements. For the user adaptive model with 22.7 min of training data, the user adaptive model with 3.78 min of training data, and the user-independent model with 0 min of user data, respectively, the defined movements have MPJPEs of 5.9, 6.8, and 8.3 cm; the dance movements have MPJPEs of 8.2, 9.5, and 10.8cm; freestyle movements have MPJPEs of 10.8, 12.1, and 13.5cm.
@@ 6.4. Results Analysis based on Motion
We further break down the results into motion types. Like many other data-driven wearable MoCap systems, the performance decreased for estimating unseen movements. As we discussed in Sec. 5, we cannot exhaust all possible upper-body movements&poses for evaluation. However, when we designed the study, we explicitly included dance and freestyle movements to better understand our system’s limitations on extreme and unseen poses and movement patterns. As shown in Fig. 10, across all three models, defined movements (5.9cm, 6.8cm, and 8.3cm with 22.7min, 3.8min, and 0min of user data) have the smallest errors, followed by dance movements (8.2cm, 9.5cm, and 10.8cm with 22.7min, 3.8min, and 0min of user data) and freestyle movements (10.8cm, 12.1cm, and 13.5cm with 22.7min, 3.8min, and 0min of user data). Adding user-specific training data also consistently improves the tracking accuracy for each movement type. This result suggested that the system similar to many data-driven systems, can further benefit from being trained with a much larger dataset containing a large variety of body movements. However, as a proof of concept, it still showed reliable tracking performance on unseen movements and subjects.
@@ 6.5. Seam Removal Ablation Study
[…]
[]This figure shows 3 bar graphs that display the Seam Removal Ablation Study results of: a User-Adaptive Model with 22.68 minutes of User Data, a User-Adaptive Model with 3.78 minutes of User Data, and a User-Independent Model with 0 minutes of User Data.The y-axis is the Mean Per Joint Position Error. The bars are color coded with gray being no seam removals, purple being sleeve removal, beige being shoulder front removal, green being shoulder back removal, and blue being shoulder top removal. For the user adaptive model with 22.7 min of training data, the user adaptive model with 3.78 min of training data, and the user-independent model with 0 min of user data, respectively, without any removals, the MPJPEs are 6.0, 7.0, and 8.6cm; with sleeve remove, the MPJPEs are 7.2, 8.0, and 8.9cm; with shoulder front removal, the MPJPEs are 7.1, 7.8, and 9.6cm; with shoulder back removal, the MPJPEs are 7.0, 8.0, and 9.7cm; with shoulder top removal, the MPJPEs are 6.3, 7.2, and 9.1 cm.
As we discussed in Sec. 4.3, the shoulder top seams have the smallest changes in magnitude, removing them does not harm the performance greatly: 0.2cm, 0.2cm, and 0.5cm worse with 22.7min, 3.8min, and 0min of user data. The shoulder front/back seams have similar significance across all models, and removing them increases the errors by 1.0cm/0.9cm, 0.8cm/1.0cm, and 1.0cm/1.1cm with 22.7min, 3.8min, and 0min of user data. On the other hand, the sleeve electrodes’ importance varies across models. For both adaptive models, the sleeve electrodes have the most important placements (i.e., their removal causes the largest errors compared with other placements) which lead to about 1cm additional error when removed. For the user-independent model, the sleeve electrodes become the least important placement and only lead to an additional error of 0.3cm. The observation implies that signals from some placements (sleeve) generalize across users worse than others do. Furthermore, given the reasonable performance across all three models with the sleeve electrodes removed, SeamPose has the potential to generalize on short-sleeve shirts and even sleeveless ones.
6.6. Performance Comparison with Prior Works
@@ 4 IMUs
[…]
[]This table situate SeamPose’s tracking performance in the literature. The table as 4 columns: the first column lists the project names; the second column lists the tracking device(s) in use; the third column lists the user-independent MPJPE in cm; the fourth column lists the user-dependent MPJPE in cm. MI-Poser, published in 2023, uses magnetic and inertial sensor fusion on a headset and 2 controllers and achieved a user-independent MPJPE of 6.6cm. PoseSonic, published in 2023, uses acoustic sensing on smartglasses and achieved a user-independent MPJPE of 6.2cm and a user-dependent MPJPE of 5.6cm. LIP, published in 2024, uses 4 IMUs on a jacket and achieved a user-independent MPJPE of 10.58cm. MoCaPose, published in 20243, uses 16 conductive patches on a jacket and achieved a user-independent MPJPE of 8.8cm and a user-dependent MPJPE of 8.6cm. SeamPose, our work, uses 8 conductive seams on a shirt and achieved a user-independent MPJPE of 8.6cm and a user-dependent MPJPE of 6.0cm.